OpenStudy (anonymous):

which is greater x>1 quant a quant b x(x2)4 (x3)3 (the numbers are exponet)

OpenStudy (anonymous):

quant b

OpenStudy (anonymous):

try pluggin in 2 and ull see

OpenStudy (anonymous):

can you demonstrate

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok but im gettin 512 for both when i plugg in 2

OpenStudy (anonymous):

x(x^2)^2 x^2=2^2=4 x(4)^2 (8) is that correct